1. Find the general solution of the system whose augmented matrix is given below. 1 -33 - 1 - 3 3 o
Answer #1

solution given that [: 3 با ما : با ما 9-33 = 3 -y-33=3 From quation we get 3+33 Putting the value of y in equation We we get